OpenStudy (anonymous):

ln a + ln x/3= 5, how do I solve for x?

OpenStudy (anonymous):

presumably your answer will have an "a" in it yes?

OpenStudy (anonymous):

is it \[\ln(x)+\ln(\frac{x}{3})=5\] or \[\ln(a)+\ln(\frac{x}{3})=5\] ?

OpenStudy (anonymous):

if it is the second one then write \[\ln(\frac{ax}{3})=5\] so \[\frac{ax}{3}=e^5\] and then \[x=\frac{3e^5}{a}\]